OpenStudy (anonymous):

Which is 256 written as a power of a number?

OpenStudy (anonymous):

2^8

OpenStudy (anonymous):

Or 4^4 :)

OpenStudy (anonymous):

256^1

OpenStudy (anonymous):

Do you have a multiple choice question? How are guesses?

OpenStudy (anonymous):

16^2 = 256, also 16^2 = 4^2 so 256 = 4^(2^)^2 = 4^4. Similarly 2^8 = 256 since it is 2^(2x2x2) you can even express it as 2^(2^3)

OpenStudy (anonymous):

sorry i mean 16 = 4^2 NOT 16^2 = 4^2

OpenStudy (anonymous):

a. 16^2 b 17^2 c 64^4 d 32^8